------------------------------------------------------------
Date: Thu Nov 15 00:10:35 2001, Shift: 16:00-24:00
Supervisor: TST, with: 
Log book#10 pages 140 - 141
Shift Summary:
------------------------------------------------------------
1:  A long store with high rates from 17.44 to 23.57
Started at ZDC rate ~ 1700 and fell slowly to ~ 150.
2:  Run      Trig6
    5798      57k      Data, MRS@40,4B    FS@12,1/2A
    5799               Pedestal run
    5800               TPC pedestal run
    5801               DVM pedestal run
    5802      55k      Data, MRS@40,4B    FS@12,1/2A
    5803      72k      Data, FS@12,1/2A   
                       DO NOT USE FOR MRS (changing polarity)
    5804     105k      Data, MRS@40,4A, FS@12,1/2A
    5805     106k      Data, same settings
    5806     114k      Data, same settings
    5807       6k      Data, same settings
Fulfilled program at MRS@40,4A and B, and have now half a
million trig6 for FS@12,1/2A - very fruitful shift (I hope!) :-)
3: No new detector problems.
------------------------------------------------------------
This archive was generated by hypermail 2b30 : Thu Nov 15 2001 - 00:11:23 EST